# Preprint no. NJU-INP 107-25 Light-Front Transverse Nucleon Charge and Magnetisation Densities

### v2

@@ I Introduction
Herein, we calculate and compare transverse charge densities obtained using both the $3$-body [27] and $q(qq)$ [29] approaches to nucleon structure. Sections II and III provide a detailed discussion of the contemporary $3$-body framework for calculation of nucleon form factors. This scheme was employed in Ref. [27], but little in the way of explanation was provided. A complete exposition of the modern $q(qq)$ approach was provided in Ref. [29], so no recapitulation is necessary. Transverse charge and magnetisation densities, and the kindred charge density associated with a transversely polarised nucleon, are defined, and the CSM predictions reported and discussed in Sec. IV. Section V provides a summary and perspective.
@@ II.1 Basic elements
$S(k;\zeta)=Z(k^{2};\zeta)/[i\cdot k+M(k^{2})]\,$, (4)
@@ IV.1 Transverse Dirac charge densities
Figure 4 depicts transverse charge densities for the proton and neutron. The primary theory curves, labelled “3-body”, were obtained from the form factors reported in Ref. [27], which were calculated using the formalism described in Sects. II, III. These curves have an associated uncertainty. It expresses a $1\sigma$-band deriving from the approach used to produce form factors at large-$\Delta^{2}$, namely, the Schlessinger point method (SPM) [74, 75, 76]. Details are provided in Ref. [27]: evidently, the impact on our results of this uncertainty is practically negligible. The second set of theory predictions was obtained using the quark + interacting-diquark, $q(qq)$, picture of nucleon structure described in Ref. [29]. Those form factors are reported as functions of $x=Q^{2}/m_{N}^{2}$. To obtain results in GeV2, we employed $m_{N}=0.94\,$GeV.

### v3

@@ I Introduction
Herein, we calculate and compare transverse charge densities obtained using both the $3$-body [27] and $q(qq)$ [29] approaches to nucleon structure. Sections II and III provide a detailed discussion of the contemporary $3$-body framework for calculation of nucleon form factors. This scheme was employed in Ref. [27], but little in the way of explanation was provided. An exhaustive background to the $q(qq)$ approach is provided in Ref. [28] and a complete exposition of the modern implementation was provided in Ref. [29], so little recapitulation is necessary. Transverse charge and magnetisation densities, and the kindred charge density associated with a transversely polarised nucleon, are defined, and the CSM predictions reported and discussed in Sec. IV. Section V provides a summary and perspective.
@@ II.1 Basic elements
In formulating and solving the relevant equations, we use a mass-independent (chiral-limit) momentum-subtraction renormalisation scheme [55], with renormalisation scale $\zeta=\zeta_{19}=19\,$GeV, viz. a far ultraviolet value so as to avoid truncation ambiguity and ensure that the the quark wave function renormalisation constant is essentially unity.
@@
$S(k;\zeta)=Z(k^{2};\zeta)/[i\gamma\cdot k+M(k^{2})]\,$, (4)
@@ IV.1 Transverse Dirac charge densities
Figure 4 depicts transverse charge densities for the proton and neutron. The primary theory curves, labelled “3-body”, were obtained from the form factors reported in Ref. [27], which were calculated using the formalism described in Sects. II, III. These curves have an associated uncertainty. It expresses a $1\sigma$-band deriving from the approach used to produce form factors at large-$\Delta^{2}$, namely, the Schlessinger point method (SPM) [74, 75, 76]. Details are provided in Ref. [27]: evidently, the impact on our results of this uncertainty is practically negligible.
The second set of theory predictions was obtained using the quark + interacting-diquark, $q(qq)$, picture of nucleon structure described in Ref. [29]. As first revealed, independently, in Refs. [77, 78, 79], the $q(qq)$ picture emerges when one exploits the pairing tendency of fermions to simplify the Faddeev equation sketched in Fig. 2. One thereby arrives at a bound-state equation that describes dressed quarks and fully-interacting quark-quark (diquark) correlations built therefrom, which bind together into a baryon, at least in part, because of the continual exchange of roles between the spectator and diquark-participant quarks. Many effectual studies of nucleon properties have employed this $q(qq)$ scheme; see, e.g., Refs. [28, 80, 23, 73, 49, 81, 82, 83, 29].
A significant merit of the dynamical $q(qq)$ approach is that it greatly simplifies the form of the Poincaré-covariant amplitude required to describe the nucleon: instead of the 128 independent tensor structures needed in the fully $3$-body treatment, only $8$ are required in the $q(qq)$ approximation. Hence, it is much easier to use and may deliver additional insights. The limitation of the Ref. [29] $q(qq)$ approach is its less rigorous connection with QCD dynamics: rather than being directly calculated, QCD-constrained Ansätze are introduced for dressed-constituent propagators and interaction vertices. One aim of the comparisons presented herein is to assist in judging the fidelity of the Ansätze used.
The form factors in Ref. [29] are reported as functions of $x=Q^{2}/m_{N}^{2}$. To obtain results in GeV2, we employed $m_{N}=0.94\,$GeV.